OpenStudy (anonymous):

A model of the population, P, of carp in a lake at time t is given by the DE: dP/dt=0.25P(1-0.0004P) A census taken 10 years ago found there were 1000 carp. Estimate the current population. When I take the integral, just get that P= (.25P^2/2)-(.0001^3/3)+C. What am I doing wrong?

OpenStudy (anonymous):

Simply taking the integral of both sides with respect to the same variable will not work here. What you're doing: \[\begin{align*}\frac{dP}{dt}&=0.25P(1-0.0004P)\\\\ \int\frac{dP}{dt}dt&=\int 0.25P(1-0.0004P)~dt\end{align*}\] The right side is like saying you want to take the integral of some unknown function \(f(t)\) with respect to \(t\). However, \(P\) is a function dependent on \(t\), so you can perform this operation unless you know what \(P\) is. Instead, you have to split the differentials (the \(d~\Box\)): \[\begin{align*}\frac{dP}{dt}&=0.25P(1-0.0004P)\\\\ \frac{dP}{0.25P(1-0.0004P)}&=dt \end{align*}\] Now you can integrate both sides w.r.t. their appropriate variables: \[\int\frac{dP}{0.25P(1-0.0004P)}=\int dt\]